# SCB-Dataset: A Dataset for Detecting Student Classroom Behavior

### v3

@@ Abstract
Using deep learning methods to detect the classroom behaviors of both students and teachers is an effective way to automatically analyze classroom performance and enhance teaching effectiveness. Then, there is still a scarcity of publicly available high-quality datasets on student-teacher behaviors. Based on the SCB-Dataset3 we proposed previously, we have introduced a larger, more comprehensive, and higher-quality dataset of student-teacher classroom behaviors, known as SCB-Dataset5. Our dataset comprises 7428 images and 106830 labels across 20 classes: hand-raising, read, write, bow head, turn head, talk, guide, board writing, stand, answer, stage interaction, discuss, clap, yawn, screen, blackboard, teacher, leaning on the desk, using the phone, using the computer. We evaluated the dataset using the YOLOv7 series of algorithms We believe that SCB-Dataset5 can provide a solid foundation for future applications of artificial intelligence in education. Our SCB-Dataset5 can be downloaded at the following link: :https://github.com/Whiffe/SCB-dataset
@@ II-B Students classroom behavior detection
III SCB-Dataset5
@@
IV Experiment
IV-A Experimental Environment and Dataset
@@ IV-B Model Training
IV-C Experimental Results and Analysis

### v6

@@ Abstract
Using deep learning methods to detect the classroom behaviors of both students and teachers is an effective way to automatically analyze classroom performance and enhance teaching effectiveness. Then, there is still a scarcity of publicly available high-quality datasets on student-teacher behaviors. Based on the SCB-Dataset3 we proposed previously, we have introduced a larger, more comprehensive, and higher-quality dataset of student-teacher classroom behaviors, known as SCB-Dataset5. Our dataset comprises 7428 images and 106830 labels across 20 classes: hand-raising, read, write, bow head, turn head, talk, guide, board writing, stand, answer, stage interaction, discuss, clap, yawn, screen, blackboard, teacher, leaning on the desk, using the phone, using the computer. We evaluated the dataset using the YOLOv7 series of algorithms We believe that SCB-Dataset5 can provide a solid foundation for future applications of artificial intelligence in education. Our SCB-Dataset5 can be downloaded at the following link:https://github.com/Whiffe/SCB-dataset
@@ II-B Students classroom behavior detection
III Annotation Work
Annotation work is the most time-consuming and labor-intensive part of SCB-Dataset5, accounting for nearly 90% of the total workload. Since 2021, we have gone through the entire process, from defining input-output expectations and behavior classifications to formulating annotation rules. However, due to numerous unreasonable and immature definitions in the early stages of dataset creation, almost a year and a half of time was wasted. By the first half of 2023, we redesigned the annotation process and introduced an extensible behavior annotation method, which allows us to flexibly expand on the existing foundation, no matter how many behaviors need to be added in the future.
III-A Improvements to Annotation Tools
To meet the practical needs of annotation work, we made multiple versions of optimizations and improvements to the annotation tool VIA. The details are as follows:
VIA Original Version
The link to the original version of VIA is as follows(as shown in Fig. 3 ): https://whiffe.github.io/VIA/via_image_annotator.html
Figure 3: Original Annotation Interface of VIA
Second Version
In this version, we optimized the label display position. As shown in Fig. 4. Labels are now displayed inside the annotation boxes instead of outside. This improvement was designed for classroom scenarios where many annotation boxes are located at the top of the image. Displaying labels inside the boxes makes it more convenient for inspection and verification. We have further optimized the function of switching the display of labels. In addition to using the mouse scroll wheel, we have added the keys ”z” and ”x” to switch the display of labels. This means that users can switch the display of labels either by using the mouse scroll wheel or by pressing the keys ”z” and ”x”. This design allows users to easily switch labels even without a mouse, making the use more convenient. https://whiffe.github.io/VIA/via_image_annotatorK.html
Figure 4: VIA Second Version
Third Version
The third version further optimized the selection of annotation boxes by introducing the mouse scroll switching feature. As shown in Fig. 5. In the original version of VIA, annotation boxes could only be selected by clicking with the mouse. If the annotation box was too small (typically caused by mislabeling), it became difficult to select. This version is particularly suitable for cleaning up small boxes created by mislabeling. Additionally, when scrolling the mouse, the selected annotation box changes color, helping users identify which boxes have been selected and which have not. This feature is especially useful for images containing a large number of targets. https://whiffe.github.io/VIA/via_image_annotator2.html
Figure 5: VIA Third Version
Fourth Version
In the fourth version, as shown in Fig. 6. We optimized the display of annotation content by showing it in half-page format, which significantly improves annotation efficiency. Additionally, we introduced two new modes: Full Image Mode and Single Target Mode.
  • Full Image Mode: Displays all annotation boxes in the entire image.
  • Single Target Mode: Displays each annotated target individually. This feature is particularly suitable for dense scenarios, allowing users to check whether each annotation box is accurately drawn and aiding in behavior classification verification and analysis.
https://whiffe.github.io/VIA/via_image_annotator3.html
Figure 6: VIA Fourth Version
Fifth Version
Building on the previous version, the fifth version introduced the copy previous frame annotations feature. As shown in Fig. 7. This functionality is particularly useful for annotating consecutive frames with high similarity, significantly reducing repetitive operations, improving annotation efficiency, and further lowering labor costs. https://whiffe.github.io/VIA/via_image_annotator4.html
Figure 7: VIA Fifth Version
Final Checks with viaJson
After completing each annotation, we use the viaJson counting website to verify the annotation results. As shown in Fig. 8. This tool identifies any unclassified annotation boxes (i.e., boxes drawn but not categorized) and provides the coordinates of the annotation boxes. Additionally, it provides statistics on the number of detection boxes and annotated targets in the current file, helping us further ensure the completeness and accuracy of the annotations. https://whiffe.github.io/VIA/via_cout_labels.html
Figure 8: Annotation Review and Counting Website
IV SCB-Dataset5
@@
V Experiment
V-A Experimental Environment and Dataset
@@ V-B Model Training
V-C Experimental Results and Analysis
@@ VI Conclusion
VII Acknowledgements
For the development of this dataset, we would like to thank the following individuals, companies and universities for their support:
  • ELEMENT TECH({CJK}UTF8gbsn成都元素科技有限公司https://www.elementech.net/)
  • SICHUAN UNIVERSITY OF SCIENCE AND TECHNOLOGY, Wang Wengang({CJK}UTF8gbsn四川科技职业技术学院,王文刚)
  • Urban Vocational College Of Sichuan, Li Jianlong({CJK}UTF8gbsn四川城市职业学院 李健龙)
  • Chengdu Neusoft University({CJK}UTF8gbsn成都东软学院)
  • Sichuan Normal University({CJK}UTF8gbsn四川师范大学
  • Beijing Normal University({CJK}UTF8gbsn北京师范大学)